# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

### v1

@@ Abstract
We introduce a new benchmark, ChartMimic, aimed at assessing the visually-grounded code generation capabilities of large multimodal models (LMMs). ChartMimic utilizes information-intensive visual charts and textual instructions as inputs, requiring LMMs to generate the corresponding code for chart rendering. ChartMimic includes $1,000$ human-curated (figure, instruction, code) triplets, which represent the authentic chart use cases found in scientific papers across various domains (e.g., Physics, Computer Science, Economics, etc). These charts span $18$ regular types and $4$ advanced types, diversifying into $191$ subcategories. Furthermore, we propose multi-level evaluation metrics to provide an automatic and thorough assessment of the output code and the rendered charts. Unlike existing code generation benchmarks, ChartMimic places emphasis on evaluating LMMs’ capacity to harmonize a blend of cognitive capabilities, encompassing visual understanding, code generation, and cross-modal reasoning. The evaluation of $3$ proprietary models and $11$ open-weight models highlights the substantial challenges posed by ChartMimic. Even the advanced GPT-4V, Claude-3-opus only achieve an average score of $73.2$ and $53.7$, respectively, indicating significant room for improvement. We anticipate that ChartMimic will inspire the development of LMMs, advancing the pursuit of artificial general intelligence.
@@ 1 Introduction
Code generation [1] is a rather demanding task that requires advanced abstract thinking and logical reasoning, reflecting the unique intelligence of human beings. Recently, progresses in artificial general intelligence (AGI) have demonstrated the potential of large foundation models [2, 3] in solving these tasks that were once the exclusive domain of human abilities [4, 5]. However, existing code generation benchmarks [6, 7, 8, 9] solely use text as input, while humans receive information from multiple modalities when coding [10, 11], such real-life scenario has not been fully explored.
@@
To this end, we present ChartMimic (Fig. 2), a multimodal code generation benchmark. ChartMimic is characterized by its (1) information-intensive visual inputs, (2) diverse chart types and (3) multi-level evaluation metrics. Specifically, compared to natural images, scientific charts convey nuanced semantic meanings through intricate visual logic, thereby exhibiting higher information density. Based on this, we define two tasks, Direct Mimic and Customized Mimic (Sec. 2.1), utilizing charts and textual instructions as inputs. These tasks challenge LMMs to generate the corresponding code for a given chart or to incorporate new data specified in the instructions, respectively. Through the collection of academic documents and scientific papers, we identify $22$ commonly used chart types and $191$ subcategories. Subsequently, we manually annotate a total of $1,000$ (figure, instruction, code) triplets for these types (Sec. 2.2). Furthermore, we establish automatic evaluation metrics from both high-level and low-level perspectives to thoroughly assess the performance of LMMs (Sec. 2.4).
We conduct examination of $14$ LMMs on ChartMimic (Sec. 3.2), including $3$ proprietary models and $11$ open-weight models. We observe that while several open-weight models can match the performance of proprietary models such as GPT-4V on public leaderboards [12], a significant performance gap still persists on ChartMimic. Specifically, the best open-weight model, Phi-3-Vision, significantly lags behind GPT-4V, with only half the performance on two tasks, indicating substantial room for improvement. Additionally, correlation analysis (Sec. 4.3) demonstrates the high correlation between our automatic metrics and human judgment. Further error analysis (Sec. 4.4) reveals that hallucination notably hinders LMMs’ performance on ChartMimic, involving the insertion of non-existent text in the ground-truth figures and confusion of similar chart types.
@@
We envision ChartMimic as a suite of benchmarks designed to guide researchers in understanding the capabilities of their LMMs. By providing a comprehensive evaluation framework, ChartMimic aims to facilitate the growth of foundation models for the community, offering insights into LMM’s various aspects, such as visual understanding, code generation, and cross-modal reasoning.
@@ 2 The ChartMimic Benchmark
In this section, we first introduce the definition of two tasks involved in ChartMimic, and then delineate the data curation process. Subsequently, we conduct a quantitative analysis to demonstrate the quality and diversity of ChartMimic and compare it with existing related benchmarks.
@@ 2.1 Task Definition
The capability of LMMs to generate code that can accurately render a specific chart demonstrates their proficiency in visual understanding, code generation and cross-modal reasoning. Specifically, given the chart $X$ and the instruction $I$, the LMM $f$ is expected to output the code $C$ that satisfies the requirements outlined in the instruction:
@@
Figure 2: The pipeline of ChartMimic. We provide $1,000$ human-curated (figure, instruction, code) triplets. We use ChartMimic to evaluate LMMs’ proficiency in the multimodal chart-to-code generation, resulting in both high-level and low-level evaluation results.
Direct Mimic. The LMM is requested to directly generate code that can reproduce the provided chart. This task evaluates the LMM’s ability to comprehensively understand and reason about the chart to generate executable code.
Customized Mimic. With additional customized data provided in the instruction, this task requires the LMM to generate code that can render a new chart encoding the customized data while maintaining the aesthetic and design of the provided chart. This task requires the LMM to understand both the chart and customized data to integrate them effectively.
@@ 2.2 Data Curation Process
ChartMimic distinguishes itself through $4$ fundamental considerations: (1) diversity of chart types, (2) balance of charts complexity, (3) reduction of data leakage, (4) integration of authentic user requirements. We keep on these principles to complete the data curation for ChartMimic through a four-step pipeline. We provide an overview here and more details in Appendix A.
@@
General Filtering. We scrape figures from source files of publications on arXiv [^1] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024. We extract figures in PDF format, yielding approximately $174,100$ figures across various domains (e.g., Physics, Computer Science, Economics, etc). We then filter the figures based on the criteria of how designers select inspiring visualization examples, including clarity and visual appeals [13, 14], color schemes [13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ figures.
Diversity and Information Density Filtering. We first initialize a data pool, creating a taxonomy of chart types and subcategories and comparing each chart against the most similar one within the corresponding category in the data pool. Should significant differences in visual complexity or information density [13] be found, the chart is added to the pool. Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication. Ultimately, this approach yields a curated collection of $279$ charts from the generally filtered $15,800$ ones.
@@
Manual Selection and Filling. In addition to sourcing from arXiv, we also collect charts from various platforms, including the Matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are selected for their unique styles not represented in our arXiv curated $279$ charts. To mitigate the risk of data leakage, we rigorously process the data and color styling of these charts, replacing existing color schemes with those unseen in our data pool while maintaining their aesthetic appeal. Consequently, we obtain $500$ prototype charts for ChartMimic.
@@
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected 500 prototype charts. Initially, the annotators are tasked with meticulously reproducing the 500 prototype charts using Python code, resulting in a set of 500 (figure, code, instruction) triplets for the Direct Mimic task. Subsequently, to simulate the scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains. They are then required to modify the corresponding codes and instructions, leading to the 500 (figure, code, instruction) triplets for the Customized Mimic task. Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
@@
Table 1: Statistics of ChartMimic. Each one in Combination, HR, Muiltidiff, and PIP is an independent subcategory. We measure code length in terms of tokens, utilizing the Llama3 tokenizer [15]. In the level count, “A/B/C” denotes the number of chart for “easy/medium/hard” level, respectively.
@@
Code Length (AVG.) 689.3 805.7 651.9 664.6 659.9 617.2 507.3 772.2 668.7 385.1 523.7 912.0
Code Length (STD.) 289.0 272.2 146.1 303.0 248.9 287.6 161.2 136.4 168.9 123.1 228.9 298.7
Level Count 53 / 40 / 7 74 / 6 / 0 17 / 13 / 0 0 / 19 / 11 15 / 10 / 0 20 / 5 / 0 13 / 7 / 0 13 / 7 / 0 2 / 8 / 5 10 / 5 / 0 7 / 3 / 0 4 / 6 / 0
@@
Count 5 5 5 5 5 5 30 25 25 10 500
Subcategories 2 3 4 4 4 4 30 25 25 10 101+(90)
Code Length (AVG.) 832.8 352.8 431.8 310.8 845.0 302.2 641.9 689.9 788.1 958.9 682.9
Code Length (STD.) 177.5 136.7 71.9 61.7 607.0 34.7 159.0 301.5 331.0 285.9 285.1
Level Count 4 / 1 / 0 0 / 0 / 5 3 / 2 / 0 5 / 0 / 0 0 / 2 / 3 5 / 0 / 0 3 / 19 / 8 1 / 4 / 20 0 / 12 / 13 0 / 0 / 10 249 / 169 / 82
2.3 Data Statistics
Notably, the (figure,code,instruction) triplets for both Direct and Customized Mimic tasks share the same figure. Therefore, we detail the data statistics for the $500$ triplets in the Direct Mimic task. As depicted on the left side of Fig. 2, ChartMimic encompasses a total of $22$ categories, with $18$ types of regular charts and $4$ types of advanced charts. For the $18$ regular chart types, we identify $101$ subcategories, with the definitions and examples of each subcategory provided in Appendix B. The advanced chart types such as Plot-in-Plot (PIP), Multidiff, and Combination are distinct forms of amalgamating multiple chart sets into a singular chart. Given the variety of their internal combination elements, each of these can be treated as a unique subcategory. Meanwhile, the Hard-to-Recognize (HR) category encapsulates unclassifiable charts, with each chart being considered a category unto itself. When we factor in the additional $90$ subcategories represented by these advanced chart types, ChartMimic encompasses a total of $191$ subcategories. This extensive diversity underscores the comprehensive nature of our benchmark. Additionally, we employ Llama3 [15] tokenizer to measure the code length. The results are shown in Table 1, which shows an average code token length of $635.98$ with a standard deviation of $253.06$. Moreover, we conduct a manual complexity assessment for each chart, classifying them into three levels: easy, medium, and hard. The distribution of charts according to complexity levels is as follows: $249$ for easy, $169$ for medium, and $82$ for hard.
@@ 2.4 Evaluation Metrics
For tasks within ChartMimic, an appropriate evaluation necessitates comparing the visual similarity between the generated and ground-truth figures. To achieve automatic evaluation, we propose multi-level metrics (i.e., low-level and high-level), which assess the similarity at different granularities. We compute the average score between the low-level and high-level scores, which range from 0 to 100, as the overall score. The illustration of the multi-level metrics is depicted in Fig. 2.
GPT-4V Score. Following the successful utilization of large foundation models like GPT-4 for automated evaluation in both natural language processing [16, 17, 18] and computer vision [19, 20], we adopt the GPT-4V score as our high-level metric. Specifically, we input both the ground-truth chart and the generated chart to GPT-4V, and instruct GPT-4V to output a high-level similarity score between 00 and $100$. The detailed prompt template can be found in Appendix C.
@@
ChartQA [21] Human Curated 3 10K I+NL NL Accuracy
Chart-to-Text [22] Crawl 6 44K I+NL NL Match-based
ChartSumm [23] Human Curated 3 84K I+NL NL Match-based
@@
HumanEval [6] Human Curated - 164 Code Code Pass Rate
MBPP [7] Human Curated - 500 NL+Code Code Pass Rate
MMCode [24] Crawl - 263 I+NL Code Pass Rate
MatPlotBench [20] Human Curated 13 100 NL Code GPT-4V Score
Design2Code [25] Crawl HTML 484 I+NL Code Multi-Level
ChartMimic (Ours) Human Curated 22 1K I+NL Code Multi-Level
In addition to high-level similarity, evaluating low-level elements between generated and ground-truth figures can help better analyze the behaviors of LMMs. Therefore, we propose evaluating four key low-level elements: text, layout, type and color. Extracting them from figures is a challenging task, as existing extraction models often fall short in terms of accuracy [26]. Considering that figures are rendered based on the code, we design a code tracer to monitor the execution process of the ground-truth code and generated code. The code tracer records the text, layout, type and color information. We calculate the F1 score of these elements as their corresponding score. And we average text, layout, type and color scores to get the low-level score. The methodology for obtaining these elements is briefly introduced as follows, with detailed descriptions provided in Appendix C.

### v2

@@ Abstract
We introduce a new benchmark, ChartMimic, aimed at assessing the visually grounded code generation capabilities of large multimodal models (LMMs). ChartMimic utilizes information-intensive visual charts and textual instructions as inputs, requiring LMMs to generate the corresponding code for chart rendering. ChartMimic includes $4,800$ human-curated (figure, instruction, code) triplets, which represent the authentic chart use cases found in scientific papers across various domains (e.g., Physics, Computer Science, Economics, etc). These charts span $18$ regular types and $4$ advanced types, diversifying into $201$ subcategories. Furthermore, we propose multi-level evaluation metrics to provide an automatic and thorough assessment of the output code and the rendered charts. Unlike existing code generation benchmarks, ChartMimic places emphasis on evaluating LMMs’ capacity to harmonize a blend of cognitive capabilities, encompassing visual understanding, code generation, and cross-modal reasoning. The evaluation of $3$ proprietary models and $14$ open-weight models highlights the substantial challenges posed by ChartMimic. Even the advanced GPT-4o, InternVL2-Llama3-76B only achieved an average score across Direct Mimic and Customized Mimic tasks of $82.2$ and $61.6$, respectively, indicating significant room for improvement. We anticipate that ChartMimic will inspire the development of LMMs, advancing the pursuit of artificial general intelligence.
@@ 1 Introduction
Code generation (Sun et al., 2024) is a rather demanding task that requires advanced abstract thinking and logical reasoning, reflecting the unique intelligence of human beings. Recently, advances in artificial general intelligence (AGI) have demonstrated the potential of large foundation models (Google, 2023; OpenAI, 2024; Anthropic, 2024; AI@Meta, 2024) to solve the tasks that are once the exclusive domain of human abilities (Achiam et al., 2023; Zhu et al., 2024). However, existing code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Lai et al., 2023) solely use text as input, while humans receive information from multiple modalities when coding (Liang et al., 2023; Fan et al., 2024). Such real-life scenarios have yet to be fully explored.
@@
To this end, we present ChartMimic (Fig. 2), a multimodal code generation benchmark. ChartMimic is characterized by its (1) information-intensive visual inputs, (2) diverse chart types, and (3) multi-level evaluation metrics. Specifically, compared to natural images, scientific charts convey nuanced semantic meanings through intricate visual logic, thereby exhibiting higher information density. Based on this, we define two tasks, Direct Mimic and Customized Mimic (Sec. 2.1), which utilize charts and textual instructions as inputs. These tasks challenge LMMs to generate the corresponding code for a given chart or to incorporate new data specified in the instructions, respectively. Through the collection of academic documents and scientific papers, we identify $22$ commonly used chart types and $201$ subcategories. Subsequently, we manually annotate a total of $4,800$ (figure, instruction, code) triplets for these types (Sec. 2.2). Furthermore, we establish automatic evaluation metrics from both high-level and low-level perspectives to thoroughly assess the performance of LMMs (Sec. 2.4).
We conduct the examination of $17$ LMMs on ChartMimic (Sec. 3.2), including $3$ proprietary models and $14$ open-weight models across parameter sizes from 2.2B to 76.0B. We observe that while several open-weight models can match the performance of proprietary models such as GPT-4o on public leaderboards (OpenCompass, 2023), a significant performance gap still persists on ChartMimic. Specifically, the best open-weight model, InternVL2-Llama3-76B, lags behind GPT-4o, with an average score gap of $20.6$ on two tasks, indicating substantial room for improvement in the open-source community. Our analysis of prompting methods (Sec. 4.2) reveals that GPT-4o can improve itself through self-reflection, which is a key manifestation of System 2 reasoning (Sloman, 1996; Kumar et al., 2024). This discovery highlights the vital role that the System 2 reasoning plays in LMMs when tackling the complex challenges presented by ChartMimic. Meanwhile, Correlation analysis (Sec. 4.3) demonstrates a high correlation between our automatic metrics and human evaluation, validating the effectiveness of these metrics. Further error analysis (Sec. 4.4) reveals that hallucinations notably hinder the performance of LLMs on ChartMimic, as they lead to the insertion of non-existent text into ground-truth figures and confusion between similar types of charts.
@@
We envision ChartMimic as a comprehensive suite of benchmarks designed to guide researchers in understanding their LMMs’ capabilities. By providing a comprehensive evaluation framework, ChartMimic aims to facilitate the growth of foundation models for the community, offering insights into various aspects such as visual understanding, code generation, and cross-modal reasoning.
@@ 2 The ChartMimic Benchmark
In this section, we first introduce the definition of two tasks involved in ChartMimic (Sec. 2.1), and then delineate the data curation process (Sec. 2.2). Subsequently, we conduct a quantitative analysis to assess the quality and diversity of ChartMimic (Sec. 2.3), establish evaluation metrics (Sec. 2.4), and compare it with existing related benchmarks (Sec. 2.5).
@@ 2.1 Task Definition
LMMs’ ability to generate chart-rendering code demonstrates their visual understanding, coding, and cross-modal reasoning skills. Specifically, given the chart $X$ and the instruction $I$, the LMM $f$ is expected to output the code $C$ that satisfies the requirements outlined in the instruction:
@@
Figure 2: The pipeline of ChartMimic. We provide $4,800$ human-curated (figure, instruction, code) triplets. We use ChartMimic to evaluate LMMs’ proficiency in the multimodal chart-to-code generation, resulting in both high-level and low-level evaluation results.
Direct Mimic. The LMMs are tasked to directly generate code that can reproduce the provided chart, thereby assessing their visual comprehension and reasoning capabilities.
Customized Mimic. The LMMs are requested to generate code for a new chart that incorporates customized data provided in the instruction while preserving the original chart’s aesthetic and design, assessing their ability to integrate visual and textual information.
@@ 2.2 Data Curation Process
ChartMimic distinguishes itself through $4$ fundamental considerations: (1) diversity of chart types, (2) balance of chart complexity, (3) reduction of data leakage, (4) integration of authentic user requirements. We keep on these four principles to complete the data curation for ChartMimic through a five-step pipeline. Here, we provide an overview here and more details in Appendix B.
@@
General Filtering. We scrape figures from source files of publications on arXiv [^1] https://arxiv.org/ that hold a CC BY 4.0 license with a publication date after February 2024 with PDF format. This yields approximately $174,100$ figures across various domains (e.g., Physics, Computer Science, Economics, etc). We then filter the figures based on the criteria of how designers select inspiring visualization examples, including clarity and visual appeal, color schemes (Bako et al., 2022; Quispel et al., 2018), and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ figures.
Diversity and Information Density Filtering. The filtering process involves two stages. In the first stage, we establish a data pool and categorize chart types. Charts with significant differences in complexity or information density are added to ensure diversity and effective communication (Bako et al., 2022). In the second stage, five experts from various fields independently evaluate the data, creating separate selection pools. We preserve the intersection of their selections and finalize the set through a voting process. This meticulous approach refines our collection to $279$ charts.
@@
Manual Selection and Filling. In addition to sourcing from arXiv, we also collect charts from various platforms, including the Matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are selected for their unique styles not represented in our arXiv curated $279$ charts. To mitigate the risk of data leakage, we rigorously process the data and color styling of these charts, replacing existing color schemes with those unseen in our data pool while maintaining their aesthetic appeal. Consequently, we obtain $600$ prototype charts for ChartMimic.
@@
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected $600$ prototype charts. Initially, annotators are tasked with meticulously reproducing the $600$ prototype charts using Python code, resulting in a set of $600$ (figure, code, instruction) triplets for the Direct Mimic task. Although other coding languages such as JavaScript and R can be used to create charts, current LMMs perform poorly when doing chart-to-code using these other languages (Sun et al., 2024). Therefore, we focus on Python code generation in the current version of ChartMimic. Subsequently, to simulate the scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains. They are then required to modify the corresponding code and instructions, leading to the $600$ (figure, code, instruction) triplets for the Customized Mimic task. Consequently, we establish the ChartMimic benchmark, comprising $1,200$ high-quality seed data.
@@
Table 1: Statistics of ChartMimic. We measure code length in terms of tokens, utilizing the Llama3 tokenizer. In the level count, “A/B/C” denotes the number of “easy/medium/hard” level, respectively.
@@
Code Length (AVG.) 689.6 794.0 681.2 685.8 689.0 655.0 529.6 779.8 655.4 418.4 624.3 975.6
Code Length (STD.) 237.8 244.4 144.7 258.7 228.2 253.0 147.1 144.3 241.4 99.5 197.7 252.3
Level Count 176 / 120 / 24 256 / 24 / 0 68 / 52 / 0 0 / 76 / 44 60 / 40 / 0 80 / 20 / 0 52 / 28 / 0 52 / 28 / 0 8 / 48 / 24 52 / 28 / 0 44 / 28 / 8 32 / 44 / 4
@@
Count 80 80 80 80 80 80 120 100 100 80 2400
Subcategories 2 3 4 4 4 4 30 25 25 20 201(101 + 100)
Code Length (AVG.) 774.4 489.4 540.0 564.5 893.4 342.2 697.4 718.9 798.2 1083.9 696.0
Code Length (STD.) 161.8 87.8 104.7 117.5 631.0 36.3 163.6 265.5 271.2 290.1 278.4
Level Count 52 / 28 / 0 0 / 28 / 52 44 / 32 / 4 56 / 24 / 0 0 / 52 / 28 52 / 28 / 0 12 / 76 / 32 4 / 16 / 80 0 / 48 / 52 0 / 0 / 80 1100 / 868 / 432
Data Augmentation. Following the development of seed triplets, we initiate a process of manual data augmentation. Annotators are tasked with altering various elements of each seed triplet, including data, color schemes, mark styles, etc., to produce augmented triplets. For each seed triplet, we create $3$ additional augmented triplets. This process enhances our dataset, yielding a total of $4,800$ triplets that reflect a wide range of realistic and practical chart use cases.
2.3 Data Statistics and Division
The (figure, code, instruction) triplets for both Direct and Customized Mimic tasks share the same figure. Therefore, we detail the data statistics for the $2,400$ triplets in the Direct Mimic task. As depicted on the left side of Fig. 2, ChartMimic encompasses a total of $22$ categories, with $18$ types of regular charts and $4$ types of advanced charts. For the $18$ regular chart types, we identify $101$ subcategories, with the definitions and examples of each subcategory provided in Appendix D. The advanced chart types such as Plot-in-Plot (PIP), Multidiff, and Combination are distinct forms of amalgamating multiple chart sets into a singular chart. Given the variety of their internal combination elements, each of these can be treated as a unique subcategory. Meanwhile, the Hard-to-Recognize (HR) category encapsulates unclassifiable charts, with each chart in the seed data being considered a category unto itself. When we factor in the additional $100$ subcategories represented by these advanced chart types, ChartMimic encompasses a total of $201$ subcategories. This extensive diversity underscores the comprehensive nature of our benchmark. We employ Llama3 (AI@Meta, 2024) tokenizer to measure the code length. As shown in Tab. 1, ChartMimic has an average code token length of $696.0$ with a standard deviation of $278.4$. Additionally, we manually categorize charts into three complexity levels: easy (1,100), medium (868), and hard (432). The detailed categorization criteria and assessment methodology are thoroughly documented in Appendix B.
We further divide the $4,800$ examples of ChartMimic into two subsets: test and testmini set. The test set comprises $3,600$ examples, while the testmini set is composed of $1,200$ examples. The testmini set is designed for rapid model development validation. Our partitioning strategy ensures each chart type is proportionally represented, preserving a distribution in the testmini set that closely aligns with the test set. Detailed comparative experimental results, discussed in Appendix C, demonstrate the consistency across two subsets. Unless otherwise stated, we report results on the testmini set.
@@ 2.4 Evaluation Metrics
For tasks within ChartMimic, an appropriate evaluation necessitates comparing the visual similarity between the generated and ground-truth figures. To achieve this, we propose multi-level metrics (i.e., high-level and low-level) to assess the similarity at different granularities. Specifically, the high-level metric encompasses the GPT-4o Score, and the low-level metric encompasses Text, Layout, Type and Color Score. We compute the average score between the high-level and low-level scores, ranging from 0 to 100, as the overall score. The illustration of the multi-level metrics is depicted in Fig. 2.
GPT-4o Score. Following the successful use of large foundation models for evaluation in both natural language processing (Zheng et al., 2024; Li et al., 2023; Dubois et al., 2024) and computer vision (Zhang et al., 2023; Yang et al., 2024; Wu et al., 2024b), we adopt GPT-4o score as our high-level metric. Specifically, we input both the ground-truth figure and the generated figure into GPT-4o, and instruct it to output a high-level similarity score ranging 00 to $100$. Although CLIP Score (Radford et al., 2021) is widely used for assessing image similarity, in our preliminary experiments, it has struggled to distinguish variations in types and other critical elements in charts, resulting in a low correlation with human evaluation results. Therefore, we use only GPT-4o Score as our high-level evaluation metric. The detailed description of GPT-4o Score can be found in Appendix E.
@@
ChartQA (Masry et al., 2022) Human Curated 3 10K I+NL NL Accuracy
Chart-to-Text (Kantharaj et al., 2022) Crawl 6 44K I+NL NL Match-based
ChartSumm (Rahman et al., 2023) Human Curated 3 84K I+NL NL Match-based
CharArXiv (Wang et al., 2024b) Human Curated 18 93K I+NL NL GPT-4 Score
@@
HumanEval (Chen et al., 2021) Human Curated - 164 Code Code Pass Rate
MBPP (Austin et al., 2021) Human Curated - 500 NL+Code Code Pass Rate
MMCode (Li et al., 2024b) Crawl - 263 I+NL Code Pass Rate
MatPlotBench (Yang et al., 2024) Human Curated 13 100 NL Code GPT-4 Score
Plot2Code (Wu et al., 2024a) Crawl 15 132 I+NL Code Multi-Level
Design2Code (Si et al., 2024) Crawl HTML 484 I+NL Code Multi-Level
ChartMimic (Ours) Human Curated 22 4,800 I+NL Code Multi-Level
In addition to high-level similarity, evaluating the similarity among low-level elements between generated and ground-truth figures can provide a more fine-grained analysis of LMMs. Therefore, we propose to evaluate four key low-level elements in charts (Savva et al., 2011; Poco & Heer, 2017): text, layout, type and color. Extracting them from figures is a challenging task, as existing extraction models often fall short in terms of accuracy (Meng et al., 2024). Considering that figures are rendered based on the code, we design a code tracer to monitor the execution processes of the ground-truth code and generated code. The code tracer records the text, layout, type and color information. We calculate the F1 score of these elements as their corresponding score. And we average text, layout, type and color scores to get the low-level score. The methodology for obtaining these elements is briefly introduced as follows, with detailed descriptions provided in Appendix E.
@@
It is important to note that code execution success rate is a standard metric for code generation tasks (Sun et al., 2024). We have implicitly incorporated this aspect into our high-level and low-level scores. Specifically, if the code fails to execute successfully, both the low-level and high-level scores are assigned a value of 0. Therefore, we do not separately weight it into the overall score.